### v2

# The Percept-V Challenge: Can Multimodal LLMs Crack Simple Perception Problems?

Abstract
Cognitive science research treats visual perception, the ability to understand and make sense of a visual input, as one of the early developmental signs of intelligence. Its TVPS-4 framework categorizes and tests human perception into seven skills such as visual discrimination, and form constancy. Do Multimodal Large Language Models (MLLMs) match up to humans in basic perception? Even though there are many benchmarks that evaluate MLLMs on advanced reasoning and knowledge skills, there is limited research that focuses evaluation on simple perception. In response, we introduce Percept-V, a dataset containing 6000 program-generated uncontaminated images divided into 30 domains, where each domain tests one or more TVPS-4 skills. Our focus is on perception, so we make our domains quite simple and the reasoning and knowledge required for solving them are minimal. Since modern-day MLLMs can solve much more complex tasks, our a-priori expectation is that they will solve these domains very easily. Contrary to our belief, our experiments show a weak performance of SoTA proprietary and open-source MLLMs compared to very high human performance on Percept-V. We find that as number of objects in the image increases, performance goes down rather fast. Our experiments also identify the perception skills that are considerably harder for all models.
The Percept-V Challenge: Can Multimodal LLMs Crack Simple Perception Problems?
Samrajnee Ghosh  and Naman Agarwal  and Hemanshu Garg  and Chinmay Mittal and Mausam and Parag Singla
1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated high performance on complex tasks, leading some researchers to suggest that these models are approaching human-level intelligence (Zhou et al., 2023; King, 2023) or even achieving MENSA-level capabilities (Schregel, 2020). However, current visual benchmarks typically assess image understanding alongside reasoning and specialized domain knowledge in fields such as engineering, medicine, and art (Nguyen, 2023; Yue et al., 2024). This integrated approach makes it challenging to isolate MLLMs’ pure image understanding capabilities from their reasoning abilities and domain-specific knowledge. To our knowledge, there is little work that focuses specifically on visual perception, the fundamental ability to understand and interpret visual images.
Visual perception represents a foundational cognitive ability that develops early in human cognition, preceding the acquisition of complex reasoning and knowledge-based skills (Rabindran and Madanagopal, 2020). Recent research has raised concerns about MLLMs’ perceptual capabilities (Zhang et al., 2024b; Fu et al., 2024), highlighting the need for a dedicated perception benchmark. To address this gap, we introduce Percept-V, a perception-focused dataset comprising 6000 program-generated, uncontaminated image-based problems from 30 domains, which primarily evaluate visual understanding while requiring minimal problem-solving skills and no specialized domain knowledge. Sample questions are presented in Figure 1.
Visual Closure - These are two images of some sequence of shapes separated by black horizontal lines - The first image consists of a sequence of shapes and the second image consists of a sequence of outlines of shapes - Determine the number of shapes in the first image that match the outlines in the corresponding place in the second image Refer to caption correct ans: 5 GPT-4o ans: “ANSWER: 3” Visual Sequential Memory - This image contains some concentric circles in different colors. - Write down the sequence of colors seen from inside to outside - The colors belong too the following list - [black,gray,brown,maroon,red,coral,tan,orange, ivory,goldenrod,yellow,green,olive,turquoise, skyblue,lavender, purple, pink, fuschsia] - No color is repeated. Refer to caption correct ans: ivory,coral,pink,green,orange,black,fucshia, yellow,skyblue,lavender GPT-4o ans: ivory,coral,pink,green,orange,black,pink, yellow, fucshia,skyblue,lavender Visual Discrimination - These are two images containing shapes such as circles, triangles and squares - The images are identical except for the fact that the second image has a few objects missing - Count the number of missing objects in the secnd image Refer to caption correct ans: 10 GPT-4o ans: “COUNT: 5” Visual Figure Ground - This is an image with some multi-coloured shapes drawn on top of it - Identify the shapes drawn over the background image - Form a set of all the unique shapes seen in the image Refer to caption correct ans: circle,pentagon,rectangle,triangle GPT-4o ans: “circle,triangle” Visual Memory - This is an image with some shapes drawn and numbers written on top of the shapes - Determine the list of numbers written on top of the circles. Refer to caption correct ans: 1,3,6,7,9,10,14,15,16,20 GPT-4o ans: “CIRCLES:1,3,7,10,14,15” Visual Spatial Relationship - There will be an image containing a rectangle and some circles above and below it - Count the number of circles above the rectangle and below the rectangle Refer to caption correct ans: “ABOVE:10,BELOW:10” GPT-4o ans: “ABOVE:8,BELOW:10” Visual Form Constancy - This is an image containing some black circles - Count the total number of circles in the image Refer to caption correct ans: 10 GPT-4o ans: “CIRCLES:10”
Figure 1: Sample questions from the Percept-V dataset illustrating tasks related to different visual perception skills as defined by the TVPS-4 framework. Each example shows the image prompt, the correct answer, and a typical incorrect response from an MLLM.
Our dataset construction draws from cognitive science literature, specifically utilizing the skill classification framework from TVPS-4 (Test of Visual Perceptual Skills) (Martin and Gardner, 2006), which is widely used for human perception assessment. TVPS-4 organizes visual perceptual abilities into seven distinct categories: visual discrimination, visual memory, visual sequential memory, visual-spatial relationships, figure-ground perception, form constancy, and visual closure (detailed in Section 3).
Skill Domain
Visual Discrimination (D) change colour, vanishing objects, comparing size, counting shapes, sort lines, sort circles, locate circles colour, locate circles shape, match shadow, match outline
Visual Memory (M) list colours, list shapes, circle boxes, colours present, count coloured circles, numbered shapes, graph counting, sort circles, sort lines, counting shapes, identifying shapes
Visual Sequential Memory (SM) grid path, layered colours, layered shapes, comparing size, match outline, match shadow
Visual Figure Ground (FG) list colours, list shapes, locate circles colour, locate circles shape, layered shapes, layered colours
Visual Form Constancy (FC) mirror image, water image, counting circles, counting shapes, numbered shapes
Visual Closure (C) match outline, match shadow, layered shapes, layered colours
Visual Spatial Relationship (SR) comparing size, circle location, circle right triangle, counting locations, cross and knots, inside circles, maze solving
Table 1: Domains using each skill. Note that a single domain can appear in multiple rows in the above table since multiple skills may be required to solve the its Problem Instances.
Percept-V comprises 30 domains, each designed to evaluate a combination of 1-3 TVPS-4 skill categories. Every domain features a single standardized question prompt that describes the specific task requirements. To systematically assess performance across varying levels of difficulty, each domain contains 200 images with varying problem size, which is operationalised through parameters such as the number of objects present, grid dimensions, or the number of processing steps required to complete the task. This structured approach enables us to evaluate MLLM performance as a function of visual complexity and provides insight into how they handle increasing perceptual demands.
We experiment on three state-of-the-art proprietary MLLMs – GPT-4o, O4-mini and Gemini 2.5 Flash and two open source MLLMs – Qwen 2.5 VL Instruct and DeepSeek VL2 Tiny, which include both language and reasoning models, and also perform a human study. Our findings reveal substantial gaps between MLLM and human performance on these simple visual perception tasks. Across all evaluated models, we observe consistent performance degradation as problem size increases, suggesting systematic limitations in handling visual complexity. While some models demonstrate relative strength in specific skill categories, overall performance on a skill remains remarkably similar across models, indicating broad deficits in visual understanding. Overall, our work is an important step towards understanding the visual perception limitations that must be addressed in the future to develop more robust MLLMs.
2 Related Work
Most exist benchmarks for evaluating the performance of MLLMs necessitate a combination of knowledge, reasoning and perception for good performance. The NTSEBench dataset (Pandya et al., 2024) comprises questions from the commonsense reasoning test from India. The MathVista dataset (Lu et al., 2023) examines different SoTA models on fine-grained, deep visual understanding and compositional reasoning. The MathVision benchmark (Wang et al., 2024) tests MLLMs on questions with a visual context with 16 distinct mathematical disciplines and 5 levels of difficulty.
OlympiadBench(He et al., 2024) collects advanced reasoning questions from the fields of mathematics, physics and chemistry to assess the complex reasoning abilities of MLLMs. Works like MMMU (Yue et al., 2024) evaluate MLLMs on many disciplines such as engineering, medicine, humanities, science, business and art. An extension of MMMU in Chinese called CMMMU (Zhang et al., 2024a) additionally tests the language understanding skills of MLLMs. SciBench (Wang et al., 2023) contains college-level science questions. Some datasets require interpretation of abstract figures, geometry shapes and scientific plots of ArXiv papers (Li et al., 2024), and experiments suggest that MLLMs could likely be used as scientific assistants in the future. EMMA Hao et al. (2025) is a recent work requiring cross-modal reasoning for answering problems in domains such as mathematics, physics, chemistry and coding. All these datasets, along with perception, require complex reasoning skills to answer questions, and several datasets additionally require specialized knowledge of a subject.
Some recent benchmarks specifically try to carefully disentangle the reasoning and perceptual abilities of LLMs. (Chung et al., 2025) propose MatHLENS which separately tests the perceptual, reasoning and integrating capabilites of MLLMs. In recent times, additional studies and benchmarks have been proposed that specifically highlight the lack of MLLMs’ perceptual abilities, such as Lee et al. (2025). Zhang et al. (2024b) shows that MLLMs perform poorly in small object and low quality object recognition. Fu et al. (2024) develops a benchmark of simple perceptual questions that can be hard for MLLMs. Most of these works use complex image scenes for evaluation and do not study the impact of problem size on perception tasks.
We construct Percept-V to fill this gap – the questions in our dataset require relatively simple reasoning and no specialized knowledge, thus better evaluating perception performance of MLLMs. Our dataset is automatically generated thus avoiding the issues of contamination, and allows for explicitly testing against the variation in problem sizes in a given domain.
Most related to our work is a very recent unpublished paper (Kanade and Ganu, 2025), which also develops a dataset in the same spirit as ours. While it is contemporaneous research, our dataset is much larger – 30 domains, compared to 12, and allows for a more fine-grained variation along problem size, making our evaluation more extensive. We ground each domain into the specific skills from TVPS-4 framework covering all 7 skills, whereas they work with only a subset of them. 43% of our domains require multiple TVPS-4 skills, which allows us to assess MLLM’s ability when applying skills in concert. Their work does not explicitly address the issue of handling multiple skills in a single domain.
Type Description # Domains
Single Boolean / Numeric 11
Fixed Length A fixed length list 6
List An ordered list 8
Set An unordered set 5
Table 2: Question prompt types based on the expected answer format, see example  1
3 Dataset
We now describe the details of our dataset referred to as Percept-V. Our dataset is composed of 30 different domains where each domain tests one or more basic perception skills from the TVPS-4 cognitive framework. The skills listed in the TVPS-4 framework are as follows:
  1. 1.
    Visual Discrimination: The ability to determine differences or similarities in objects based on size, colour, shape, etc.
  2. 2.
    Visual Memory: The ability to recall visual traits of a form or object.
  3. 3.
    Visual Sequential Memory: The ability to recall a sequence of objects in the correct order.
  4. 4.
    Visual Spatial Relationships: - Understanding the relationships of objects within the environment.
  5. 5.
    Visual Figure Ground: The ability to locate something in a busy background.
  6. 6.
    Visual Form Constancy: The ability to know that a form or shape is the same, even if it has been made smaller/larger or has been turned around.
  7. 7.
    Visual Closure: The ability to recognize a form or object when part of the picture is missing.
Each domain in our dataset consists of multiple problem instances which require a subset of TVPS-4 skills Martin and Gardner (2006) for them to be solved correctly. Table 1 lists the domains for each skill, which require the knowledge of that skill in order for its problem instances to be solved. In total, 17 domains test one skill, 8 domains have two skills, and 5 domains test 3 skills in concert.
Each domain is associated with a unique question prompt. Further, each problem instance in a domain is composed of a pair $(Q,I)$, where $Q$ represents the question prompt associated with the domain, and $I$ is an image generated automatically through Python programmes that used libraries like Matplotlib, Pillow and OpenCV.111image generation for each domain is controlled by a set of hyperparameters described in A. In other words, we apply the same generic question prompt to different images resulting in different problem instances. The problem instances are categorized across multiple sizes, where size of a problem instance captures its inherent complexity, such as number of objects or number of distractors, or number of sequence of steps required to solve the problem instance. For example, in the counting_circlescounting\_circles domain shown under Visual Form Constancy in 1, the problem size is controlled by controlling the number of circles to be counted. On the other hand, in the numbered_shapesnumbered\_shapes domain under Visual Memory in 1, with increasing problem size, the number of circles as well as distractors i.e., triangles and squares increase. Further, each question prompt, $Q$ is composed of three types of prompts $(in,r,op)$, where $in$ provided the description of the input image $I$, $r$ provided details of the task to be performed and $op$ specified the answer format of the model output.
Each domain has its problem instances divided across the sizes varying from 1 to 20, with 10 instances generated for each size in each domain in the dataset. This results in 200 problem instances for each domain and a total of 6000 problem instances divided across 30 domains. Examples of problem instances from our dataset are provided in Figure 1.
There are four types of question prompts in our dataset based on the type of the answer they expect. Details are listed in Table 2. The answer can be a single value (Boolean, numeric), a fixed length answer (a tuple or a list of a fixed length for all problem sizes), an ordered list (a list that has to maintain a specific sequence with its length dependent on the problem size) or an unordered set (a set with only unique values). For each answer type, the question prompt also lists a specific format in which the answer is to be outputted, so that it can be automatically evaluated. All MLLMs mostly honor the format consistently – the errors are from incorrect perception and not format mismatches.
We note all the images in our dataset are generated using automated programs. This has multiple advantages: (1) it helps control the complexity of the problem instances created, and (2) it helps avoid any potential contamination issues, which has been pointed out as a serious concern in some of the existing benchmarks (Shojaee et al., 2025). Further details of the images in each domain along with specific examples, as well as our automated generative process, is described in the Appendix.
4 Experiments
Through our experiments, we try to answer the following questions:
  1. 1.
    How well do the SOTA open as well as proprietary LLMs perform on Percept-V?
  2. 2.
    Do some LLMs perform better than others especially for a subset of skills?
  3. 3.
    How does the LLM performance vary across variation in skills: are some skills harder than others?
  4. 4.
    How does the LLM performance vary with increasing size of the problem instances in different skills?
  5. 5.
    Finally, how does the LLM performance compare with human performance on Percept-V?
In order to answer these questions, we perform our experimental comparison using three different state-of-the-art proprietary LLMs: (1) GPT-4o (2) Gemini 2.5 Flash (3) o4-mini, and two open source MLLMs: (1) Qwen-2.5-7B-Instruct (2) Deepseek-VL2-Tiny. We note that among these, o4-mini and Gemini 2.5 Flash are reasoning/thinking models. For open source models, all our experiments were run on an NVidia A100 GPU. More details about model parameters and GPU runtime are presented in A. The hyperparameters like temperature and top_p are both kept as 0.0 to produce more deterministic reproducible answers from models like Gemini, GPT-4o and Qwen (for the remaining models, the temperature could not be made 0.0 due to model restrictions). We run all experiments in a zero-shot mode to keep the costs under control; further, our preliminary experiments on a subset of models suggest that few shot prompting does not yield any significant gains for our domains.
For each problem size of each domain, all 10 candidate problems are evaluated on all models and the average performance on these 10 problems is considered as the model accuracy for that problem size. For each LLM for a given domain, we use exactly the same prompt. This prompt is constructed by appending the question prompt followed by details of output format. Detailed examples of prompts are given in the Appendix.
Skill / Domain GPT-4o o4-mini Gemini Qwen DeepSeek Avg. Acc.
— Performance by Skill —
Visual Discrimination 23.70 55.30 51.25 11.50 4.25 29.20
Visual Memory 33.55 47.55 46.55 19.59 1.50 29.75
Visual Sequential Memory 21.50 47.50 41.92 4.67 5.42 24.20
Visual Figure Ground 25.42 63.25 62.83 7.58 0.08 31.83
Visual Form Constancy 27.30 48.90 50.30 25.60 2.50 30.92
Visual Closure 18.40 40.20 34.60 4.60 6.50 20.86
Visual Spatial Relationship 50.36 62.64 69.79 27.79 25.21 47.16
Average of skills 28.60 52.19 51.03 14.48 6.49 37.04
— Performance by Domain —
change colour 18.50 32.00 23.50 8.00 4.50 17.30
circle boxes 33.00 34.00 28.00 12.50 10.50 23.60
circle location 41.50 46.00 57.50 14.50 6.00 33.10
circle right triangle 82.00 99.50 99.00 25.00 50.00 71.10
colours present 1.50 4.50 4.50 0.00 0.00 2.10
comparing size 36.50 31.00 55.00 5.00 0.00 25.50
count coloured circles 44.00 37.50 44.50 39.00 5.00 34.00
counting circles 57.50 55.00 55.50 42.50 4.00 42.90
counting locations 20.50 36.50 33.50 15.50 0.00 21.20
counting shapes 44.00 58.50 61.00 41.00 0.00 40.90
cross and knots 57.50 100.00 96.00 23.50 47.50 64.90
graph counting 31.00 33.50 36.00 26.00 0.00 25.30
grid path 0.50 53.00 23.50 0.00 0.00 15.40
identifying shapes 69.00 92.50 47.00 41.50 0.00 50.00
inside circles 87.50 95.50 93.00 96.50 73.00 89.10
layered colours 25.00 33.50 34.00 0.50 0.00 18.60
layered shapes 11.50 20.00 18.50 9.50 0.00 11.90
list colours 18.00 35.50 58.00 1.50 0.50 22.70
list shapes 70.50 90.50 92.00 21.00 0.00 54.80
locate circles colour 11.50 100.00 82.00 6.00 0.00 39.90
locate circles shape 16.00 100.00 92.50 7.00 0.00 43.10
match outline 23.00 84.50 72.50 7.50 15.00 40.50
match shadow 32.50 63.00 48.00 5.50 17.50 33.30
maze solving 27.00 30.00 54.50 14.50 0.00 25.20
mirror image 12.50 29.00 24.00 21.50 5.00 18.40
numbered shapes 12.00 75.50 88.50 8.00 0.00 36.80
sort circles 24.00 34.00 30.00 20.00 0.50 21.70
sort lines 22.00 27.00 22.50 5.00 0.00 15.30
vanishing objects 9.00 23.00 25.50 10.00 5.00 14.50
water image 10.50 26.50 22.50 15.00 3.50 15.60
Full Percept-V dataset 31.65 52.70 50.75 18.10 8.25 32.29
Table 3: Comparison of LLM performance across different skills and domains. All values are percentages.
4.1 Main Results
This section answers our main experimental question (Q1, Q2 and Q3). Table 3 presents the performance of various LLMs across different skills (aggregated over domains that test that skill), as well as over each individual domain. There are several observations to make.
Domain-wise Analysis:
For a significant majority of the domains, across LLMs the performance is below par. In general, as expected, closed source models do better than open source. Among closed source models, GPT-4o which is a non-thinking model, does the worst, with its performance on more than 12 (out of a total of 30) domains 20% or less, and on 24 domains 50% or less. Thinking models do somewhat better, with these numbers being 2 (out of 30), and 17, respectively, for o4-mini, which is closely followed by Gemini. Open source models are in general below par with 20 domains performing 20% or less, and only a single domain performing better than 50%, for Qwen. DeepSeek has similar performance. In terms of overall average performance across domains as well, o4-mini is the best performing model with its accuracy only at 52.7%.
Figure 2: The overall accuracy of all models in different skills.
These results clearly point to the inability of LLMs, both closed source, and open, as well thinking and non-thinking, to be able to achieve any meaningful performance on most of the domains in our Percept-V dataset. This is contrast with humans, which perform significantly better on these tasks (see Section 4.3). We also see that LLMs are relatively consistent in their performance across domains, for example, ’inside circles’ has best overall performance, and this is also the domain, where most LLMs achieve their (near) best. Similarly, ’colors present’ has overall the worst performance, and this is also the domain, where individual LLMs also perform poorly compared to other domains. This domain requires models to count objects of a given color, and we hypothesize that MLLMs get confused in related colors (such as turqouise and blue). Overall, this suggests that some domains are inherently harder than others for all LLMs.
Skill-wise Analysis:
A skill-wise analysis of performance presents no different story. o4-mini is best performing with average accuracy across skills at 52.19% closely followed by Gemini. [^2] We note that average accuracy across skills may not match average across across domains, since the same domain may be repeated across multiple skills. The best performing o4-mini can achieve more than 50% accuracy only on 3 out of a total of 7 skills, with its higher accuracy on any skill being 63%, demonstrating that it struggles on majority of the skills to get a reasonable performance. GPT4o sees a sharp decline in performance with average accuracy at 28.6%, and among open source models, Qwen is only at 14.48%. We see that o4-mini and Gemini are consistently better than GPT4o across skills, followed by open source models, which perform quite poorly. Similar to the case of domains, there is consistent behavior of LLM performance skill-wise, with some skills being harder (easier) than others for all LLMs.
For instance, most LLMs perform their best on Visual Spatial Relationship, and (near) worst on Visual Closure. Visual spatial relationships generally test relations such as an object being to the left of or above another object. Since image captions generally contain such relationships, it is likely that MLLMs got trained well on this skill and hence perform well. On the other hand, visual closure questions test on inside-out relationships, layered objects, or matching a solid object with just its outline. Such phenomena are likely absent from multimodal training data, making models weaker on this skill.
4.2 Performance vs Size
In this section, we set out to answer Q4: what is the variance in LLM performance across varying size? Results comparing the skill-wise model performance across different problem sizes are presented in Figure 2. There are 7 graphs, one for each skill, and the last graph in the bottom row presents performance of each LLM averaged over skills. We observe that performance of models drop significantly with the increase in problem size for all skills. We observe that for most of the skills, thinking models (o4-mini and Gemini) solve smaller problems with lesser number of objects sufficiently well, with near-perfect accuracy, for several skills. However, this performance drops dramatically with increasing size. As seen in Table 3, o4-mini and Gemini are among the best performing models, and are relatively robust to size change compared to other LLMs, despite a performance drop. These are followed by GPT-4o, and open source models perform the worst. Clearly, these results indicate that LLMs are not able to generalize across size across different skills. We hypothesize that the key reason for this performance drop is their inability to ’discern’ the underlying pattern of the task to be solved in a given domain, resulting in sharp decline in performance as the size increases. It is also possible that for thinking models, they are able to ’discern’ this pattern, but are unable to operationalize this for larger sizes.
4.3 Human Study
Figure 3: Comparison of MLLM average accuracy versus human accuracy across the seven TVPS-4 skills.
This section answers Q5, i.e., how does LLM performance vary in contrast with human subjects. A study on 16 college students was done with the Percept-V. It had 6 female and 10 male subjects. The subjects were presented randomly selected questions of any problem size from the 30 domains. Results of Human Study are presented in Figure 3. Here, for fair comparison, the models are also evaluated on the exactly same samples as were given to the humans. It can be seen that models lag severely as compared to humans. Even the best performance of the models exhibited by Gemini in Visual Spatial Relationship is around 60% as compared to about 90% accuracy shown by humans in the same skill. Overall see Table 4 – human performance is about 30 percent points higher than than closest MLLM. When testing on ‘colours present’, the domain on which all MLLMs have abysmal performance, we find the gap to be humongous – humans achieve 81.25% performance whereas best MLLMs on the same questions achieve only 6.25%.
This experiment underscores a large gap in the overall perceptual ability of MLLMs compared to humans – an important direction for future work.
Model Overall performance
GPT-4o 31.17%
o4-mini 51.46%
Gemini 49.27%
Qwen 20.40%
Deepseek 8.45%
Human 81.67%
Table 4: Overall performance in human study
4.4 Compliance with Output Format
We find that most proprietary models hardly falter in following the answer formats specified in the question, and most of their accuracy issues stem from lack of perceptual skills alone. However, this is not true for DeepSeek, which is the worst at following output formats in all categories. The best performing is GPT-4o (with o4-mini close second) that sticks to the output formats with 0% format errors. Single answer format is the easiest to follow for models, whereas fixed answer type is the hardest as can be seen from Table 5.
Answer Type GPT-4o o4-mini Gemini Qwen DeepSeek
single 0.00 0.00 0.00 0.00 3.68
fixed 0.00 0.17 0.83 5.18 61.58
list 0.00 0.00 0.00 0.00 22.44
set 0.00 0.10 0.20 19.10 60.00
Table 5: Format Error (%) Analysis of models
5 Conclusion
Our paper takes an important step forward in assessing the perceptual abilities of multimodal LLMs. To isolate perceptual ability, our questions use very few and basic reasoning skills (such as counting), and similarly very basic knowledge (of colors and basic shapes). We present Percept-V, an extensive dataset of 6000 automatically generated images from 30 domains, which are uncontaminated, i.e., not seen by any LLM before. We also annotate each domain with their perceptual skills from the TVPS-4 Cognitive Science framework.
When tested on our dataset, all MLLMs perform rather weakly on almost all domains. The best average performance on our dataset is only 52.7% (o4-mini), and open models perform rather poorly (8-18%). As complexity of questions increases, all model performances degrade quite fast on almost all skills, with thinking models continuing to perform decently on a couple of skills.
We also perform human study to compare human performance with MLLMs. We find that the gap between the best MLLM and humans is quite high (over 30 percent points), suggesting that contrary to LLMs’ high performance against humans in many tasks (e.g., LLMs winning Math olympiad), MLLMs are rather behind, especially on basic perception skills. We will release Percept-V upon publication. We hope that it will become a standard dataset for testing any new MLLM, and will help in improving the quality of future MLLMs.
Limitations
This paper evaluates MLLMs only against a limited pool of adult college students. Future work may be expanded into including annotators from different age groups and diverse educational background to improve the robustness of the results. Additionally, this paper only identifies the weakness of MLLMs in visual perceptual skills and does not investigate in detail, the rationale behind such weakness.
Ethical Considerations
All human participants were shown the instructions required to be followed during the experiment, and they volunteered to participate in the study. Additionally, all humans were asked for consent and financially compensated for their contribution – details can be found in Appendix A. Our work only tries to expose the perceptual limitations of SoTA MLLMs, so human study did not elicit any personally identifying information – the participants were simply answering simple perception questions.
Appendix A Appendix
A.1 Domain examples of MLLMs
A.1.1 change color
Prompt:
-These are two identical images with circles of different colours. - The colours of the circles at the same position in the two images may be different. - Count the number of differently coloured circles between the two images. - The output must be given in a single line in the form of COUNT:x
A.1.2 circle boxes
Prompt:
- This is an image with some red circles divided and a blue partition that divides the image into two sides. - Some circles are on the left side of the partition whereas some are on the right. Note that, a side of the image may be completely empty as well. - Determine the number of circles to be moved from the side with more number of circles to the side with less number of circles so that each side has equal number of circles. - In case of both the sides having equal number of circles, return zero. - The output must be given in a single line in the form of COUNT:x where x is the number of circles to be transferred
A.1.3 circle_location
Prompt:
- This is an image with some black circles drawn in the four quadrants - The quadrants are numbered in the default way i.e. quadrant 1 is the rightmost and topmost quadrant, quadrant 2 is the leftmost and topmost, quadrant 3 is the leftmost and lowest and quadrant 4 is the rightmost and lowest. - Determine the quadrant with the most number of black circles and the count of black circles in that quadrant. - In case of a tie, return the lesser quadrant number. - The output must be given in a single line in the form of QUADRANT:x COUNT:y where x is the quadrant number and y is the count of circles.
A.1.4 circle_right_triangle
Prompt:
- This is a grid which contains some circles and triangles - Check if any triangle cell has a circle cell to its immediate right. - The last line of output should be "NO" if no such cell exists and "YES" otherwise.
A.1.5 colours_present
Prompt:
- This image contains some shapes in different colours. - The colours belong to the following list - [black, gray, brown, maroon, red, coral, tan, orange, ivory, goldenrod, yellow, green, olive, turquoise, skyblue, blue, lavender, purple, pink, fuchsia] - No colour is repeated - Produce a list of yes or no; write yes in the list if the colour at the corresponding index of the above list is present and no if it is absent.
- The output should be written in a single line as a comma-separated list of yes or no, written in order of the given list of colours. - For example, the output should be of the form ANSWER: yes , no, no, yes and so on.
A.1.6 Comparing Size
Input Prompt:
- This is an image with several rows of circles where each row is above a horizontal line.
- Each row contains two colored circles (one blue and one green) of different sizes.
- Analyse which circle is bigger between the two in each row.
- The last line of the output should be of form "ANSWER:" followed by the space separated color answers for all the rows in the image.
- For example ANSWER: Blue Green Blue.
A.1.7 Count Colored Circles
Input Prompt:
- This is an image containing some colored circles.
- This is an image containing some colored circles.
- Last line of the output must be of the form COUNT:x
A.1.8 Counting Circles
Input Prompt:
- This is an image containing some black circles.
- Count the total number of circles in the image.
- The last line of the output should be of form COUNT:x.
A.1.9 Counting Locations
Input Prompt:
- There will be an image containing a rectangle and some circles above and below it.
- Count the number of circles above the rectangle and below the rectangle.
- The output must be of form in a single line ABOVE:x BELOW:y
A.1.10 Counting Shapes
Input Prompt:
- This is an image containing shapes such as circles, triangles and squares.
- Count the number of circles, triangles and sqaures in the image.
- The last line of the output must be of the form CIRCLES:x TRIANGLES:y SQUARES:z
A.1.11 Cross and Knots
Input Prompt:
- This is a grid containing some black X.
- Each cell is given a coordinate (x,y) which is (row,column) with 0 based indexing.
- An adjacent cell to a cell A is any cell sharing an edge with A
- A safe cell is a cell that is not adjacent to any black X.
- Output the coordinates of any one safe cell as labelled in the grid.
- If no such cell exists output None.
- The last line of the output should just contain the cell coordinates (x,y) or None and nothing else.
A.1.12 Graph Counting
Input Prompt:
- This is a graph where the blue colored circles are nodes and the lines joining them are edges.
- Count the number of nodes and edges in the graph.
- Last line of the output must be of the form NODES:x EDGES:y.
A.1.13 Grid Path
Input Prompt:
- This is a grid which contains some circles, triangles and squares.
- Travel from cell labelled S to cell labelled E.
- Follow the path marked by the black line and arrows.
- Write down the sequence of shapes seen while moving along this path.
- The last line of the output should be a comma-separated list of shapes without colors, written in order of visiting of form SHAPES : Shape 1 , Shape 2 and so on.
A.1.14 Identifying Shapes
Input Prompt:
- This is an image with several rows of different shapes where each row is above a horizontal line. - Each row contains a coloured shape.
- Identify the shape present in each row.
- The ouput should have only a single line containing the answer.
- The output format should be "ANSWER:" followed by the space seperated shape names.
- For example, ANSWER: Circle Triangle Circle
A.1.15 Inside Circles
Input Prompt:
- This is an image containing some black circles and a single red dot.
- Count the total number of circles in the image which contain the red dot
- Note that a circle containing the red dot means the centre of the dot is present inside the circumference of the circle
- The last line of the output should be of form COUNT:x.
A.1.16 Layered Colors
Input Prompt:
- This image contains some concentric circles in different colours.
- Write down the sequence of colours seen from inside to outside.
- The colours belong to the following list - [black, gray, brown, maroon, red, coral, tan, orange, ivory, goldenrod, yellow, green, olive, turquoise, skyblue, blue, lavender, purple, pink, fuchsia]
- No colour is repeated
- The output should be written in a single line as a comma-separated list of colours, written in order from the innermost colour to the outermost colour
- For example, the output should be of the form COLOURS: Colour 1 , Colour 2 and so on.
A.1.17 Layered Shapes
Input Prompt:
- This image contains some shapes that are layered on top of each other in a white background.
- The image could also contain a single shape with no other shapes layered on top of it.
- Write down the sequence of shapes from the inside layer to the outside layer.
- If there is a single shape then identify it.
- The last line of the output should be a comma-separated list of shapes without colors, written in order from top to bottom of form SHAPES : Shape 1 , Shape 2 and so on.
- If there is a single shape then write SHAPES : shape-name where shape-name is the name of the shape seen
A.1.18 List Colors
Input Prompt:
- This is an image with some multi-coloured shapes drawn on top of it
- The colours belong to the following list - [black, gray, brown, maroon, red, coral, tan, orange, ivory, goldenrod, yellow, green, olive, turquoise, skyblue, blue, lavender, purple, pink, fuchsia]
- No colour is repeated
- Identify the colour of the shapes drawn over the background image
- Form a list of all the colours of the shapes
- The output must be given in a single line in the form of a list of all the colours found
A.1.19 List Shapes
Input Prompt:
- This is an image with some multi-coloured shapes drawn on top of it.
- Identify the shapes drawn over the background image
- Form a set of all the unique shapes seen in the image
- The output must be given in a single line in the form of a set of unique shapes found.
A.1.20 Locate Circles Color
Input Prompt:
- This is a grid which contains some coloured circles.
- The rows and columns are numbered in the grid.
- The coordinate of a cell is given by the row number followed by the column number.
- Write down all the coordinates of the cells which contain a green circle.
- The last line of the output should be a space-separated list of the coordinates of the cells which contain a green circle.
- The coordinates should be in the format (row,column) withh the comma and bracket.
A.1.21 Locate Circles Shape
Input Prompt:
- This is a grid which contains some coloured circles , rectangles and triangles.
- The rows and columns are numbered in the grid.
- The coordinate of a cell is given by the row number followed by the column number.
- Write down all the coordinates of the cells which contain a green circle.
- The last line of the output should be a space-separated list of the coordinates of the cells which contain a green circle.
- The coordinates should be in the format (row,column) with the comma and bracket.
A.1.22 Match Outline
Input Prompt:
- These are two images of some sequence of shapes separated by black horizontal lines
- The first image consists of a sequence of shapes and the second image consists of a sequence of outlines of shapes
- Determine the number of shapes in the first image that match the outlines in the corresponding place in the second image.
- The output must be given in a single line in the form of ANSWER:x where x is the number of matching shapes
A.1.23 Match Shadow
Input Prompt:
- These are two images of some sequence of shapes separated by black horizontal lines.
- The first image consists of a sequence of shapes and the second image consists of a sequence of shadows of shapes
- Determine the number of shapes in the first image that match the shadows in the corresponding place in the second image.
- The output must be given in a single line in the form of ANSWER:x where x is the number of matching shadows.
A.1.24 Maze Solving
Input Prompt:
- This is a maze. - You start from the cell labelled S.
- Your goal is to reach cell labelled E through the shortest path possible.
- You can move up, down, left or right.
- You cannot move through the black walls.
- You can only move through white cells.
- Give the sequence of cells taken to go from S to E.
- Last line of output should be a comma separated line with cell numbers starting with S and ending at E and nothing else.
subsubsectionMirror Image Input Prompt: - These are two images containing different shapes in different colours - The image consists of a vertical blue line at the centre. - Some objects of the first image are mirrored along the given vertical line in the second image. - Other objects in the second image maintain the same postions as in the first image. - Determine the number of objects that are mirrored in the second image. - The output must be given in a single line in the form of COUNT: x, where x is the number of being mirrored
A.1.25 Numbered Images
Input Prompt:
- This is an image with some shapes drawn and numbers written on top of the shapes.
- Determine the list of numbers written on top of the circles.
- The output must be given in a single line in the form of CIRCLES:x where x is the list of numbers seen on top of the circles
A.1.26 Sort Circles
Input Prompt:
- There are several black circles with different sizes.
- Sort these circles by size, from smallest to largest.
- Output must only contain the space separated labels of circles in their sorted order by size.
A.1.27 Sort Lines
Input Prompt:
- The image has sseveral parallel lines of varying lengths.
- Sort these lines by length, from shortest to longest.
- Last line of the output must only contain the space separated labels of thee lines in their sorted order by length and nothing else.
A.1.28 Vanishing Objects
Input Prompt:
- These are two images containing shapes such as circles, triangles and squares
- The images are identical except for the fact that the second image has a few objects missing
- Count the number of missing objects in the secnd image
- The output must be given in a single line in the form of COUNT:x
A.1.29 Water Image
Input Prompt:
- These are two images containing different shapes in different colours
- The image consists of a horizontal blue line at the centre.
- Some objects of the first image are mirrored along the given horizontal line in the second image.
- Other objects in the second image maintain the same postions as in the first image.
- Determine the number of objects that are mirrored along the horizontal line in the second image.
- The output must be given in a single line in the form of COUNT: x, where x is the number of being mirrored
A.2 Image Generation
Image is generated using hyperparameters like –num_image that give the number of images to be generated per problem size and –num_size that give the list of problem sizes
A.3 Financial Compensation for Participants
Each participant was paid financial remuneration based on the number of hours taken to finish the questionnaire which was typically around one hour for 30 questions.
A.4 Instructions for Participants
Thank you for participating in this experiment! This questionnaire will test your visual perception skills through some simple questions. Please follow the below instructions carefully - 1. Focus on accuracy and not on speed. There is no time bound , so you can take breaks while solving. However, you have to keep the website open on some tab/window otherwise, your progress will be lost. Also, take sufficient time to solve each question. 2. Please open this questionnaire on laptop/tablet . Some fonts are small, so, you may miss out if you open on smaller devices. Zoom in on images that appear tiny otherwise, this may affect your accuracy! 3. Each question follows a particular answer format. Follow the answer format strictly! Check for spelling mistakes while writing. 4. Try to attempt all the questions. If you do not understand any question, reach out to us! All the best!
A.5 GPU Hours
GPU Hours required were around 5 for Deepseek (1B parameters) and Qwen(7B parameters). The number of parameters for o4-mini is 8B, GPT-4o is 200B.